# Humans as a Calibration Pattern: Dynamic 3D Scene Reconstruction from Unsynchronized and Uncalibrated Videos

### v1

@@ 5       $\tilde{R}^{i}_{\text{est}}\leftarrow R^{i}_{\text{est}}R^{\top}$
7       return $\{\tilde{R}^{i}_{\text{est}},\tilde{\tau}^{i}_{\text{est}}\}$
@@ A.3 Evaluation Details
Figure 8: We demonstrate camera pose estimation results of the initialization stage at the top row (Initialization) and the final results of the joint optimization with K-Planes at the bottom row (Refinement). Red frustums are the ground-truth camera poses and blue frustums are the estimated camera poses.
@@
Appendix B Additional Results
We visualize both initialized and refined camera poses at the top row and bottom row, respectively, in all of the scenes in our dataset in Fig. 8. We can observe that our initialization step produces good initial points, and our joint optimization with dynamic NeRF produces near-perfect pose alignments across all scenes.
Furthermore, we show additional qualitative comparisons in Fig. 9. We also provide videos rendered at the test viewpoint in the supplementary material. We recommend the readers to see the videos.

### v2

@@ 5       $\tilde{R}^{i}_{\text{est}}\leftarrow R^{i}_{\text{est}}R^{\top}$
7       return $\tilde{R}^{i}_{\text{est}},\tilde{\tau}^{i}_{\text{est}}$
@@
Figure 9: Visual illustration of estimated camera poses from our initialization stage on EgoBody dataset. Red and blue frustums are the ground-truth and estimated camera poses, respectively.
Figure 10: Additional qualitative results on Mobile-Stage dataset.
@@
Appendix C Additional Results
We additionally visualize both initialized and refined camera poses in all of the scenes in EgoBody dataset in Fig. 9 and Panoptic Studio in Fig. 11. We can observe that our initialization step produces good initial points, and our joint optimization with dynamic NeRF produces near-perfect pose alignments across all scenes.
Furthermore, we show additional qualitative comparisons in Panoptic Studio dataset in Fig. 12 and Mobile-Stage dataset in Fig. 10. We also provide videos rendered at the test viewpoint in the supplementary material. We recommend the readers to see the videos.
Figure 11: We demonstrate camera pose estimation results of the initialization stage on Panoptic Studio datsaet at the top row (Initialization) and the final results of the joint optimization with K-Planes at the bottom row (Refinement). Red frustums are the ground-truth camera poses and blue frustums are the estimated camera poses.